# Rational angle bisectors on the coordinate plane and solutions of Pell’s equations

### v17

@@ Abstract
On the coordinate plane, the slopes of two straight lines and the slope of one of their angle bisectors satisfy the equation Recently, an explicit formula for nontrivial integral solutions of this equation with solutions of negative Pell’s equations was discovered by the author. In this article, for a given square-free integer and a given integer we describe every integral solution of such that are coprime by using the fundamental unit of and elements of whose absolute value of norms are the smallest prime powers. We also describe every nontrivial rational solution of the above equation as one of its applications.
@@ 1 Introduction
Problem 1 has the meaning when the bisectors of can be drawn by connecting and other lattice points for given lattice points which is important in drawing techniques. Furthermore, in engineering, we can specify the radiation range and the axis of light with a ratio of integers without error due to approximations to irrational numbers by using a solution to this problem. Problem 1 is reduced to solving the equation
@@
where is a square-free integer
@@
Let be a square-free integer and let be an integer
@@
Let be positive integral solutions of We say is smaller than if
@@
Let denote the additive monoid of all integers
Let be a square-free integer Let be the fundamental unit of the real quadratic field Let be the fundamental solution of and put
Put
For each put
@@
and put
Put
Furthermore, for each where we denote the conjugate and norm of by
@@
Case : are the -components of rational solutions of the equation Then is given by
for some such that
Case : are the -components of rational solutions of a common negative Pell’s equation for some square-free integer Then is given by
for some such that and In addition, are written of the form
@@
The result in Case (I) refines the previous result [5, Theorem 2]. For any rational solution of ( ‣ 1) in Case (I), each of and is a ratio of the leg lengths of some Pythagorean triangle. We find the rational solutions of ( ‣ 1) in Case (II) by using the following Theorem. This is the second main theorem of this article.
@@
Let be an integer
Suppose that or “ and ” or Then has a strictly primitive integral solution if and only if
(1.5) for some for each prime number where
Suppose that and Then has a strictly primitive integral solution if and only if
(1.6) for some for each prime number
In these cases, its strictly primitive integral solution satisfies
@@
(1.8) and where in the latter case.
- (2)
@@ 2 Rational solutions of Pell’s equations
In this section, we prove Theorem 3. Throughout the section, let be a square-free integer and use the symbols in Notation 1. We recall the following classical theorem.
@@
The negative Pell’s equation has a rational solution, if and only if has no prime factors congruent to modulo
@@
See [1, Theorem 3.2.1].
By Fermat’s two squares theorem (see [8, Theorem 2.15]), can be expressed as for some if and only if has no prime factors congruent to modulo We can assume since Under these conditions, is a rational solution of
Conversely, if has a rational solution with then which implies that has no prime factors congruent to modulo by Fermat’s theorem as above, and so does ∎
@@
If has an integral solution, then has no prime factors congruent to modulo however, the converse is false. For example, has a rational solution but no integral solutions.
@@
Let be a prime number, and let be integers If has a strictly primitive integral solution then defined by
@@
For each let be an integral solution of and put Suppose that the integral solution of defined by (2.1) is not strictly primitive. Then for some
Case: Then and therefore
Case: Then This implies that divides
and therefore in Hence
@@
Case: Then and therefore
Case: Then and therefore in With respect to an extension of the -adic additive valuation to this implies
@@
Case: Then and therefore
Case: Then In this implies and therefore since Hence
@@
Suppose that
If then
Both and hold if and only if In this case, holds.
If and is a multiple of then every integral solution of is a pair of even numbers, and not strictly primitive.
If or “ and ” or then
If and then
Unless “ and ” then
for each
@@
Put Recall that the integer ring of is
Let be the monoid of all ideals in and let be the equivalence relation on defined by
which is compatible with the multiplication. Since the ideal class group of is finite, and so is the quotient monoid which can be regarded as its submonoid. Furthermore, splits in if and is unramified in if
@@
To prove the converse, suppose that an integer has a prime factor which does not split in Let be an integral solution of
Case: is unramified in or “ and is unramified in ” In the principal ideal is decomposed as
@@
(2.3) by the assumption. Hence and therefore
Case: ramifies in Then and therefore divides which implies
In these cases, is not strictly primitive. Thus the assertion of (1) holds.
Suppose that Then has no integral solutions, since
This implies that if
@@
Conversely, if and then a strictly primitive integral solution of satisfies
which implies since
@@
which imply that every integral solution of is a pair of even numbers for each multiple of
If and then since has no strictly primitive integral solutions for each integer by (2). Combining this fact with (1) and (2), we obtain the desired equalities.
The desired inequality follows from the argument in the proof of (1).∎
@@
Put and let denote the integer ring of For each prime number which splits in we take a prime ideal in lying over in and put Note that are primitive for such a prime ideal in and any Let be a strictly primitive integral solution of Let be the set of all prime factors of and put For each let and be the quotient and remainder, respectively, when dividing by For each put The equality can be expressed as with ideals in which implies
@@
in by Theorem 6 (1), and therefore the ideal is divisible by either or ; otherwise, in and also by (2.3), and therefore which contradicts that are coprime. For each if or we can see that is divisible by either or by the same argument, which implies
- (i)
Case: and Then is unramified in and by Theorem 6. The prime ideal divides both and only once by its self-conjugacy; otherwise, in and therefore in since
(2.4) which contradicts that are coprime. The ideal is decomposed as
@@
Letting we obtain (1.7).
- (iii)
Case: and Then and has no strictly primitive integral solutions for each multiple of by Theorem 6, which implies The ideal is decomposed as
with in The ideal
Letting
and defining by
where the first double sign is not necessarily the same as the second, we obtain (1.7).
The exponent satisfies (1.8), since for each and
by (see [7, Theorem 2.1.4]). Thus the assertion of (1) holds.
Let be an integral solution of and let be the greatest common divisor of and Let be the set of all prime factors of and put . Put
@@
(2.5) for some satisfying (1.8) and the conditions obtained by replacing with in (1.5), (1.6) because of (1). Put
Case or . Then and since
Case . Then and since and
Case or . Then
Case . Then
Multiplying both sides of (2.5) by
@@ 3 Rational angle bisectors
The combinations of prime factors of which divide odd times coincide with each other, since are square numbers. Let be the product of these prime factors (define if there are no such prime factors). Then there exist such that
@@
Case : are the -components of rational solutions of respectively. Let
@@
Case : are the -components of rational solutions of for some square-free integer respectively. In general, if and then

### v18

@@ Abstract
On the coordinate plane, the slopes and of two straight lines and the slope of one of their angle bisectors satisfy the equation Recently, an explicit formula for nontrivial integral solutions of this equation with solutions of negative Pell’s equations was discovered by the author. In this article, for a given square-free integer and a given integer we describe every integral solution of such that and are coprime by using the fundamental unit of and elements of whose absolute value of norms are the smallest prime powers. We also describe every nontrivial rational solution of the above equation as one of its applications.
@@ 1 Introduction
Essentially, Problem 1 has the meaning when the bisectors of can be drawn by connecting and other lattice points for given lattice points which is important in drawing techniques. Furthermore, in engineering, we can specify the radiation range and the axis of light with a ratio of integers without errors due to approximations to irrational numbers by using a solution to this problem. Problem 1 is reduced to solving the equation
@@
where is a positive square-free integer.
@@
For each square-free integer such that has an integral solution, we denote the -th smallest positive integral solution of by Every nontrivial integral solution of ( ‣ 1) is given by
for some integers after switching and if necessary, where (1.1) contains the case when Conversely, every triple of Form (1.1) or (1.2) is an integral solution of ( ‣ 1).
@@
Let be a square-free integer, and let be an integer.
@@
Let be positive integral solutions of We say is smaller than if or “ and ”
@@
Let denote the additive monoid of all nonnegative integers. For each prime number let be the normalized -adic additive valuation. We denote the greatest common divisor of by
Notation 2.
Let be a square-free integer. Let be the fundamental unit of the real quadratic field Let be the fundamental solution of and let Let
For each let
@@
and let
Let
Furthermore, for each where we denote the conjugate, norm of by
@@
Suppose that are the -components of rational solutions of the equation Then is given by
(1.3) for some such that
Suppose that are the -components of rational solutions of a common negative Pell’s equation for some square-free integer Then is given by
(1.4) for some such that and In addition, can be written in the form
@@
Conversely, every triple of Form (1.3) or (1.4) is a rational solution of ( ‣ 1).
@@
The result in (i) refines the previous result [5, Theorem 2]. For any rational solution of ( ‣ 1) in (i), each of and is a ratio of the leg lengths of some Pythagorean triangle. We find the rational solutions of ( ‣ 1) in (ii) by using the following Theorem. This is the second main theorem of this article.
@@
Let be an integer. Then has a strictly primitive integral solution if and only if
(1.5) for some for each prime number where “ and ” if In this case, its strictly primitive integral solution satisfies
@@
(1.7) and
Let be an integer. Then every integral solution of satisfies
(1.8) for some where (1.7) and
(1.9) hold.
@@ 2 Rational solutions of Pell’s equations
In this section, we prove Theorem 3. Throughout the section, let be a square-free integer. We recall the following classical theorems.
Theorem 4 ([1, Theorem 3.2.1]).
@@
The negative Pell’s equation has a rational solution, if and only if has no prime divisors congruent to modulo
@@
By Fermat’s two squares theorem (see [8, Theorem 2.15]), can be expressed as for some if and only if has no prime divisors congruent to modulo We can assume since Under these conditions, is a rational solution of
Conversely, if has a rational solution with then which implies that has no prime divisors congruent to modulo by Fermat’s theorem as above, and so does ∎
@@
If has an integral solution, then has no prime divisors congruent to modulo however, the converse is false. For example, has a rational solution but no integral solutions.
@@
Let be a prime number, and let be integers. If has a strictly primitive integral solution then defined by
@@
For each let be an integral solution of and let Suppose that the integral solution of defined by (2.1) is not strictly primitive. Then for some
Case 1: Suppose that Then and therefore
Case 2: Suppose that Then In this implies that divides
and therefore Hence
@@
Case 1: Suppose that Then and therefore
Case 2: Suppose that Then and therefore in With respect to an extension of to this implies
@@
Case 1: Suppose that Then and therefore
Case 2: Suppose that Then In this implies and therefore since Hence
@@
Let Recall that the integer ring of is
Let be the monoid of all ideals in For each we define if
for some Then the relation on is an equivalence relation which is compatible with the multiplication. Since the ideal class group of is finite, and so is the quotient monoid which can be regarded as its submonoid. Furthermore, splits in if and is unramified in if
@@
To prove the converse, suppose that an integer has a prime divisor which does not split in Let be an integral solution of
Case 1: Suppose that is unramified in Then or “ and ” by assumption. In the principal ideal is decomposed as
@@
(2.4) by assumption. Hence and therefore
Case 2: Suppose that ramifies in Then and therefore divides which implies
In these cases, is not strictly primitive. Thus the desired assertion holds. ∎
Corollary 1.
Unless “ and ” then is not greater than the ideal class number of for each
Proof.
The desired assertion follows from (2.3). ∎
In the arguments below, we use the following lemma.
Lemma 1.
Suppose that and let be an integer.
If then has no integral solutions. Furthermore, if then
Both and hold if and only if In this case, holds.
If and then every integral solution of is a pair of even integers. Furthermore, if then
If and then every integral solution of is a pair of even integers.
Proof.
If then has no integral solutions, since by assumption. This implies that if
@@
Conversely, if and then a strictly primitive integral solution of satisfies which implies since
(2.5) -
Suppose that Then
which imply that every integral solution of is a pair of even integers if This implies by (2).
@@
which imply that every integral solution of is a pair of even integers if
Now we can determine the elements of
Theorem 8.
If or “ and ” or then consists of all prime numbers which split in
If and then consists of all prime numbers which split in and
Proof.
We also use the following fact.
@@
Let and let denote the integer ring of For each prime number we take a prime ideal in lying over in and let Note that are primitive for each if splits in Let be a strictly primitive integral solution of Let be the set of all prime divisors of and let For each let and be the quotient and remainder, respectively, when dividing by For each let The equality can be expressed as with ideals in which implies
@@
in by Theorem 7, and therefore the ideal is divisible by either or ; otherwise, in and also by (2.4), and therefore which contradicts that are coprime. For each we can see that is divisible by either or by the same argument as above, which implies
The latter case follows from since
(2.6) and the fundamental solution of is strictly primitive.
Case 2: Suppose that and
Assume that Then is unramified in by Theorem 8, and by Lemma 1 (1), (4). The prime ideal divides both and only once by its self-conjugacy; otherwise, in and therefore in by (2.6), which contradicts that are coprime. The ideal is decomposed as
@@
This implies and therefore which contradicts that are coprime.
Thus and therefore by Theorem 8.
In these cases, the exponent satisfies (1.7), since
by (2.5) and (see [7, Theorem 2.1.4]). We can verify that (1.5) is satisfied by the argument above. Thus the assertion of (1) holds.
Let be an integral solution of and let Let be the set of all prime divisors of and let . Let
@@
(2.7) for some satisfying (1.7) and the condition obtained by replacing with in (1.5) because of (1). If then (1.9) holds, since If then Multiplying both sides of (2.7) by
@@ 3 Rational angle bisectors
For any prime number the parities of and coincide with each other, since are square numbers. Let be the product of every prime number such that these valuations are odd (define if there are no such prime numbers). Then there exist such that
@@
Suppose that are the -components of rational solutions of respectively. Let
@@
Suppose that are the -components of rational solutions of for some square-free integer respectively. In general, if and then